E8B.6 Electrons moving perpendicular to the direction of a magnetic field are observed to circulate around the field direction at a frequency of 150 MHz. What is the magni- tude of the magnetic field?
Answer #1

we know that

qB/m = 2*pi*f

where, q is charge on particle, b is magnetic field, m is mass of particle and f is frequency

In the given question,

q=1.6-10-19 C

m= 9.11*10-31 kg

f=150*106 Hz

B=2*pi*f*m/q=(2*pi*150*106*9.11*10-31)/(1.6*10-19)

= 0.005366 T